# RationalVLA: A Rational Vision-Language-Action Model with Dual System

### v1

@@ V-B Rational Manipulation
Quantitative Results. The quantitative results are shown in Table II. The 3DDA achieves a low success rate primarily because its language encoder provides a limited language ability, rendering it incapable of comprehending unseen instructions or distinguishing defective instructions. The LLaVA + 3DDA system achieves exceptionally high first-task success rates. This stems from its well-designed prompts that guide the pretrained VLM to first assess the executability of language instructions. This design fully unleashes the low-level action policy’s manipulation capabilities, ensuring it operates without interference from defective instructions. However, some unseen instructions and defective instructions are difficult to distinguish, requiring in-depth understanding and reasoning. LLaVA occasionally confuses these instructions. GPT-4o shows substantial improvements in both multimodal perception and language reasoning capabilities, resulting in reduced susceptibility to defective instructions and higher first-task success rates. Nevertheless, for the VLM-as-identifier system, while the high-level VLM can help the low-level policy reject defective instructions, it does not directly enhance the language generalization capabilities of the system. This is because the VLM can not directly translate unseen complex instructions for policies and further execute corresponding actions. Consequently, its advantages in subsequent task success rates and average trajectory length remain marginal.
Compared with the best baseline GPT4o + 3DDA, our method RationalVLA yields an improvement from 1.30 to 2.26 in average length. The success rate of the last task is elevated by 14.5%. RationalVLA learns a <REJ> token to enable rejections for non-executable tasks, which enables rejections for non-executable tasks. Thus, the appearance is little affected by defective instructions. In the following tasks, the high-level VLM translates complex, unseen instructions into an understandable latent action embedding, which is then taken as input by the low-level policy to generate effective actions.
@@
Figure 7: Visualization of the rejection mechanism. The defective instructions are colored in burgundy. At the right of every task, we denote the success or failure. The first row shows a failure case of the fine-tuned RoboFlamingo, it fails and terminates in task 1, with a total execution time of 76s. This perturbation makes the model inefficient and incapable of completing sequential tasks. The rest of the rows show a successful case of our RationalVLA , it successfully finishes all the tasks and uses only 48s, showing the efficiency when tackling multiple tasks.
@@ V-C Manipulation with Unseen Instructions
Results. While RoboFlamingo leverages pretrained vision-language alignment, it struggles with nuanced instruction understanding and accurate task grounding under this enriched setting III. Similarly, VLM-as-Identifier architectures fail to distinguish unseen instructions, often misclassifying them or defaulting to irrelevant behaviors.
In contrast, our method, RationalVLA, combines the MLLM for instruction interpretation with a robust multitask policy. This design enables the model to semantically parse unseen instructions and accurately associate them with corresponding actions. Notably, RationalVLA achieves an average gain of 1.05 average length over the strongest baseline, significantly outperforming all comparison methods in this setting. These results demonstrate superior generalization and interpretability in enriched manipulation tasks.
@@ V-D Manipulation With Seen Instructions
Results. As shown in Table III, RationalVLA achieves the highest 48% 5/5 task success rate and an average correct step length 0.18 higher than the strongest baseline [16]. This result demonstrates that RationalVLA remains highly competitive on standard manipulation benchmarks. Importantly, it not only retains the control effectiveness of the underlying policy, but also enhances task performance by leveraging the VLM’s language understanding and multimodal reasoning through seamless integration.
@@ VI-A Robotic System Setup
We built a teleoperated robot system to collect high-quality demonstrations and conduct experiments. The system includes a lead arm and a follower arm. The follower arm consists of a 6-DOF AgileX Piper robotic arm with a 1-DOF gripper and an ORBBEC Dabai depth camera for first-view image. We also utilize a RealSense D515 camera to capture third-view images. The other Piper arm with a teaching gripper serves as the lead arm. During data collection, the lead arm is controlled by a human expert, and the follower arm mirrors the same action. For each task with different levels of difficulty, we collect over 25 demonstrations through human operation, accompanied by annotated instructions. During the test, only the follower arm is activated, and each task is tested over 20 episodes.
@@ VI-B Basic Tasks With Unseen Instructions
Figure 8: Real-world robotic system setup. We employ a 6-DOF AgileX Piper arm equipped with a 1-DOF gripper and utilize a RealSense D515 camera to capture third-view images.
@@ VI-C Long-Horizon Tasks
  • Lift-put: “Lift the watermelon” serves as the defective instructions with no watermelon on the scene. Then the model should finish “put the pepper into the bowl”.
  • Put-put: The defective instruction is “Put the grape in the cup”.However, there is no cup on the scene. The second task is “put the watermelon into the bowl”.
@@
Results. We report success rates in Table VI and trajectory visualizations in Fig 9, which demonstrate that our RationalVLA effectively generalizes to real-world manipulation tasks even under defective instructions. Unlike 3DDA, which relies solely on a policy-based architecture and often executes all instructions indiscriminately, RationalVLA learns to reject infeasible commands in the current environment, thereby preserving the integrity of subsequent tasks. This results in an average 63.7% improvement in success rate, underscoring the benefits of incorporating instruction comprehension into the control pipeline for robust real-world deployment.
@@ VII Conclusion
This paper analyzes the limitation of the classic manipulation task and proposes a new benchmark, RAMA, which allows defective instructions. It contains more than 14k data samples. To better manage RAMA benchmark, we propose RationalVLA. It is a dual-system vision-language-action model with the capability of processing complex instructions with defective ones. Extensive experiments show that our model demonstrates surprisingly effective performance in both classic manipulation tasks and rational manipulation tasks. RAMA benchmark research further highlights its value in industrial and domestic applications. We hope our work will accelerate the deployment of robot assistants in real-world applications.
@@ Appendix A Acknowledgment
We sincerely thank Yuxin Huang for creating the exquisite video. Pengxu Hou assisted with hardware setup. We are grateful to Chris McCool for his valuable discussions and suggestions.
Figure 10: Defective Instruction Generator for RationalVLA. Our defective instruction generator is divided into two parts, Modular Generator and Direct Generator. The Modular Generator is designed to produce defective instructions for the visual, physical, semantic, and motion dimensions by systematically replacing specific linguistic components (Adj., Noun., Verb.) while maintaining strict control over variables. The Direct Generator focuses on generating defective instructions of the safety and out-of-context dimensions, leveraging carefully crafted prompts to query GPT-4o.
@@ B-A Instruction Generator
The defective instruction generator used for RationalVLA  is shown in Figure 10. Our defective instruction generator is divided into two parts, Modular Generator and Direct Generator.
@@ B-B Data Examples
The RAMA dataset encompasses defective instructions across six dimensions: visual, physical, semantic, motion, safety, and out-of-context. Examples illustrating each of these dimensions are provided in Figure 11. To enhance the diversity and generalizability of RAMA benchmark, we extend the original dataset with mixed instructions that involve multiple variable perturbations.

### v2

@@ V-B Rational Manipulation
Quantitative Results. The quantitative results are shown in Table II. The 3DDA achieves a low success rate primarily because its language encoder provides a limited language understanding ability, rendering it incapable of comprehending unseen instructions or distinguishing defective instructions. The LLaVA + 3DDA system achieves exceptionally high first-task success rates. This stems from its well-designed prompts that guide the pretrained VLM to first assess the executability of language instructions. This design fully unleashes the low-level action policy’s manipulation capabilities, ensuring it operates without interference from defective instructions. However, some unseen instructions and defective instructions are difficult to distinguish, requiring in-depth understanding and reasoning. LLaVA occasionally confuses these instructions. GPT-4o shows substantial improvements in both multimodal perception and language reasoning capabilities, resulting in reduced susceptibility to defective instructions and higher first-task success rates. Nevertheless, for the VLM-as-identifier system, while the high-level VLM can help the low-level policy reject defective instructions, it does not directly enhance the language generalization capabilities of the system. This is because the VLM can not directly translate unseen complex instructions for policies and further execute corresponding actions. Consequently, its advantages in subsequent tasks’ success rates and average trajectory length remain marginal.
Compared with the best baseline GPT4o + 3DDA, our method RationalVLA yields an improvement from 1.30 to 2.26 in average length. The success rate of the last task is elevated by 14.5%. RationalVLA learns a <REJ> token to enable rejections for non-executable tasks, which enables rejections for non-executable tasks. Thus, the performance is little affected by defective instructions. In the following tasks, the high-level VLM translates complex, unseen instructions into an understandable latent action embedding, which is then taken as input by the low-level policy to generate effective actions.
@@
Figure 8: Visualization of the rejection mechanism. The defective instructions are colored in burgundy. We use the red cross and green checkmark to denote the success and failure, respectively. The first row shows a failure case of the fine-tuned RoboFlamingo. It fails and terminates in task 1, with a total execution time of 76s. This perturbation makes the model inefficient and incapable of completing sequential tasks. The rest of the rows show a successful case of our RationalVLA, it successfully finishes all the tasks and uses only 48 s, showing the efficiency when tackling multiple tasks.
@@ V-C Manipulation with Unseen Instructions
Results. While RoboFlamingo leverages pretrained vision-language alignment, it struggles with nuanced instruction understanding and accurate language grounding under unseen, complex instructions. Similarly, VLM-as-Identifier architectures fail to transfer the understanding of unseen instructions to low-level policies, resulting in irrelevant behaviors.
In contrast, our RationalVLA combines the VLM for instruction interpretation with a robust multitask policy. This design enables the model to semantically parse unseen instructions and accurately associate them with corresponding actions. Notably, RationalVLA achieves an average gain of 1.05 average length over the strongest baseline, significantly outperforming all comparison methods in this setting. These results demonstrate superior generalization and interpretability in enriched manipulation tasks.
@@ V-D Manipulation With Seen Instructions
Results. As shown in Table III, RationalVLA achieves the highest 48% success rate in the fifth task and an average length 0.18 higher than the strongest baseline [16]. This result demonstrates that RationalVLA remains highly competitive on standard manipulation benchmarks. Importantly, it not only retains the control effectiveness of the underlying policy, but also enhances task performance by leveraging the VLM’s language understanding and multimodal reasoning through seamless integration.
@@ VI-A Robotic System Setup
We built a teleoperated robot system to collect high-quality demonstrations and conduct experiments. The system includes a lead arm and a follower arm. The follower arm consists of a 6-DOF AgileX Piper robotic arm with a 1-DOF gripper and an ORBBEC Dabai depth camera for first-view image. We also utilize a RealSense L515 camera to capture third-view images. The other Piper arm with a teaching gripper serves as the lead arm. During data collection, the lead arm is controlled by a human expert, and the follower arm mirrors the same action. For each task with different levels of difficulty, we collect over 25 demonstrations through human operation, accompanied by annotated instructions. During the test, only the follower arm is activated, and each task is tested over 20 episodes.
@@ VI-B Basic Tasks With Unseen Instructions
Figure 9: Real-world robotic system setup. We employ a 6-DOF AgileX Piper arm equipped with a 1-DOF gripper and utilize a RealSense L515 camera to capture third-view images.
@@ VI-C Long-horizon Tasks
  • Lift-put: “Lift the watermelon” serves as the defective instructions with no watermelon in the scene. Then the model should finish “put the pepper into the bowl”.
  • Put-put: The defective instruction is “Put the grape in the cup”. However, there is no cup in the scene. The second task is “put the watermelon into the bowl”.
@@
Results. We report success rates in Table VI and trajectory visualizations in Fig 10, which demonstrate that our RationalVLA effectively generalizes to real-world manipulation tasks even under defective instructions. Unlike 3DDA, which relies solely on a policy-based architecture and often executes all instructions indiscriminately, our RationalVLA learns to reject infeasible commands in the current environment, thereby preserving the integrity of subsequent tasks. This results in an average 63.7% improvement in success rates, underscoring the benefits of incorporating instruction comprehension into the control pipeline for robust real-world deployment.
@@ VII Conclusion
This paper analyzes the limitation of the classic manipulation task and proposes a new benchmark, RAMA, which allows defective instructions. It contains more than 14k data samples. To better manage RAMA benchmark, we propose RationalVLA. It is a dual-system vision-language-action model with the capability of processing complex instructions with defective ones. Extensive experiments show that our model demonstrates effective performance in both classic manipulation tasks and rational manipulation tasks. RAMA benchmark research has the potential in industrial and domestic applications. We hope our work will accelerate the deployment of robot assistants in real-world applications.
@@ Appendix A Acknowledgment
We sincerely thank Yuxin Huang for creating the exquisite video. Zhide Zhong provides vaulable suggestions and helps with real-world experiments. Pengxu Hou assisted with hardware setup. We are grateful to Chris McCool for his valuable discussions and suggestions.
Figure 11: Defective Instruction Generator for RAMA. Our defective instruction generator is divided into two parts, Modular Generator and Direct Generator. The Modular Generator is designed to produce defective instructions for the visual, physical, semantic, and motion dimensions by systematically replacing specific linguistic components (Adj., Noun., Verb.) while maintaining strict control over variables. The Direct Generator focuses on generating defective instructions of the safety and out-of-context dimensions, leveraging carefully crafted prompts to query GPT-4o.
@@ B-A Instruction Generator
The defective instruction generator used for RAMA is shown in Figure 11. Our defective instruction generator is divided into two parts, Modular Generator and Direct Generator.
@@ B-B Data Examples
The dataset encompasses defective instructions across six dimensions: visual, physical, semantic, motion, safety, and out-of-context. Examples illustrating each of these dimensions are provided in Figure 12. To enhance the diversity and generalizability of RAMA benchmark, we extend the original dataset with mixed instructions that involve multiple variable perturbations.